# Dual-component stellar assembly histories in local elliptical galaxies via MUSE

### v1

@@ 4.4.2 Star formation histories from mass fractions
The 2D SFHs built from the mass fractions in general illustrate a picture where the inner component of the elliptical galaxies are extremely old and have higher metallicities. These weights are concentrated in the upper right corner of the grid, with a minute spread in age ( Gyr), while there is a relatively larger spread in metallicity for galaxies J020536 and J225546, between -0.5 and 0.3 dex. The outer component also shows a dominant old stellar population, with a relatively wider spread in age compared to the inner component, and an even greater spread in metallicity, with stellar populations spanning ages Gyr and the entire metallicity range of the grid. These are indicative of minor episodes of continuous and slower stellar mass assembly that have occurred after the primary peak of star formation, wherein different stellar populations have been accreted.
@@
The 1D mass-weighted SFHs trace the cumulative mass assembly of each component through cosmic time. They complement the results from the 2D SFHs with a clear depiction of the inner component having assembled its stellar mass initially and swiftly within a very short timescale, forming the inner core of the galaxy. The peak mass assembly of the outer component exhibits a delay of several Gyr compared to the inner component, and has assembled its mass at a slower rate, over the course of a few Gyr, most likely growing its mass by continuous accretion during this time until attaining the present-day mass. The SFHs for the inner and outer components exhibit a flat slope from Gyr and Gyr ago respectively, up to , marking their quiescence during this period. This clearly points to an inside-out mass growth scenario for local elliptical galaxies.
@@
Furthermore, the timescales and were measured from the cumulative mass fractions (Table 4). measures the lookback time when half of the present-day stellar mass had been assembled, while measures the same for 90% of the present-day mass, which is related to the onset of quenching in the galaxy (Ferré-Mateu et al., 2018; Romero-Gómez et al., 2024), especially for retired quiescent galaxies. For all three ellipticals, the inner component had already formed half of their mass by Gyr ago, while the outer component had taken Gyr longer to assemble the same (with timescales of 2.09 Gyr, 4.74 Gyr, and 670 Myr for J020536, J205050, and J225546 respectively). Similarly, the inner component appears to have quenched its star formation activity as early as Gyr ago. The outer component, on the other hand, stopped forming stars only Gyr ago, roughly Gyr after half the stellar mass had assembled. The relative contribution of the inner component to the total stellar mass was found to be 50% for all three galaxies, indicating that half of the stellar mass originates from the outer component. These collective findings demonstrate an early and rapid assembly of the majority of the stellar content within the inner component of the elliptical galaxies, which consequently appears to be a “co-dominant” component in terms of its influence on the present day total stellar mass. The longer assembly timescales attributed to the outer component, along with it hosting half of the galaxy stellar mass, imply a slower addition of stars well after the inner component has quenched its star formation.
@@ 4.4.4 Star formation histories from light fractions
The SFHs drawn from the light fractions are shown in the lower panels of Figures 7, 8, and 9 respectively. The 1D cumulative light-weighted age distributions (right-most panels of the above-mentioned figures) serve as a qualitative proxy for the SFH, reflecting the relative contributions of stars formed at different cosmic epochs, to the current galaxy luminosity. Since the light-weighting is performed with the band luminosity, the results also trace when star formation episodes had occurred during more recent times. and derived from the cumulative light fractions (Table 4) indicate the lookback time when the stars contributed to 50% and 90% respectively of the luminosity of the galaxy or its components. In this case, does not necessarily highlight the onset of quiescence, but rather highlights the epoch when the stars contributed to 90% of the present-day luminosity. We find that the light-weighted 2D SFHs follow the trends observed in the mass-weighted SFHs, revealing a similar pattern in the history of star formation. In J205050, a significant portion of the light originates from stellar populations older than 5 Gyr, spanning a broad range of metallicities from super-solar to very sub-solar, reflecting a diverse range of stellar populations, similar to the mass-weighted 2D SFH. In contrast, the outer component of J225546 is dominated primarily by stars older than 10 Gyr, but with a wide spread in metallicities.
@@
The light fractions of J205050 similarly show a predominantly old and metal-enhanced light-weighted stellar population in the inner component, which similar to J020536, has undergone a rapid event of bulk star formation at early epochs. The lookback time, indicating the major star formation in this component, seems to have occurred and ceased roughly 13.5 Gyr ago, followed by a flat unperturbed quiescent slope leading to the present-day. The outer component, however, shows a delayed contribution to 90% of the luminosity, with most of these stars having formed around 7 Gyr ago. It has remained quiescent since then, with no significant star formation up to the present day.
@@
Following this trend, the 2D light-weighted SFH of J225546 shows an inner component again dominated by the light of the old metal-rich stars at solar and super-solar metal abundances. The outer component consists of two distinct light-weighted stellar populations, one with old and metal-enhanced stars, and the other with similarly old but quite metal-poor stars reaching [M/H] dex. The cumulative light fractions complement this with a clear indication that the peak of star formation in both the inner and outer components happened at earlier lookback times. The outer component followed after the principal star formation occurred, but the formation timescale between both components for this galaxy is not as substantial as for the other two galaxies, which is also evident from their , indicating a delay of roughly 3 Gyr.
@@
The stellar population analyses and the reconstructed star formation histories support the two-phase scenario introduced in Sect. 1. The inner regions of the ellipticals ( kpc) could, in principle, be associated to the in-situ star formation within the galaxy as a result of dissipative collapse or major mergers very early on in their lifetimes. In contrast, the outer component ( kpc in this sample), is likely a combination of star formation formed in-situ from gas, with the majority of the stellar envelope having been accreted through dry mergers over longer timescales compared to the inner component. Regarding the former, Choi et al. (2024) find that in cosmological simulations, mergers can result in the accretion of gas from external galaxies. They also find signatures of smoothly accreted gas, particularly through the cooling of the galaxy halo, aligning with the findings of Lagos et al. (2015) in ETGs. In Choi et al. (2024), another origin is attributed to gas from the primary galaxy halo that has been recycled over generations of stellar evolution. While these processes were examined in the context of gas accretion in the inner regions of galaxies fuelling a supermassive black hole, a smaller fraction of the gas can be distributed over different regions, including the outskirts. On the other hand, the accretion of direct stellar material from dry mergers with external satellites of different metallicities occurs around the old, metal-rich core which undergoes minimal evolution since its star formation was quenched early on.
@@ 5.1 Multiple structural components in ellipticals: implications from surface brightness profiles
The notion that elliptical galaxies may consist of multiple components has been examined in several studies through 2D photometric decomposition, which are outlined in this Section. This work builds upon existing methods by incorporating the imaging and spectroscopic capabilities of integral field spectroscopy. The spectro-photometric decomposition technique with BUDDI not only models the surface brightness profiles of various components at each wavelength of an IFU datacube, but also enables a more accurate derivation of stellar populations from the resulting spectra compared to photometric colours, which often suffer from the age-metallicity degeneracy (Worthey, 1999).
@@
Huang et al. (2013a) delve into investigating the possibility of multi-component ellipticals in a sample 94 local galaxies in the CGS survey. The high spatial resolution of deep optical images from this survey allowed for a reliable 2D decomposition of the ellipticals into 3-4 sub-components using GALFIT (Peng et al., 2002, 2011) to study their structural properties in the band. They find that for a majority of their sample, the surface brightness profiles are best modelled by three Sérsic components, representing a compact inner core, a middle component, and an outer extended envelope. The 2D residuals consistently improved with increasing complexity in the models. Interestingly, they report only a small fraction of their sample to be modelled with two components. Furthermore, they also suggest that their high-luminosity sample of ellipticals shows photometric signatures of being “core” ellipticals with a shallow slope within in the innermost component. Lacerna et al. (2016) report similar findings, noting a good fit with three Sérsic models for a sample of 89 isolated local elliptical galaxies from SDSS images. Huang et al. (2013a) and Huang et al. (2013b) link the inner and intermediate components to the phase of elliptical galaxy formation involving in-situ dissipative processes, while associating the outer component with a later stage of ex-situ star accretion through dry mergers. These findings and interpretations closely align with those presented in this work, despite differences in the number of components modelled.
@@
Along these lines, Spavone et al. (2017) perform a detailed analysis of the surface brightness profiles of six massive ETGs in A VST Early-type GAlaxy Survey (VEGAS) survey (Capaccioli et al., 2015). They adopt various profile combinations to identify the most accurate model. Four galaxies were effectively modelled using a Sérsic + exponential profile, whereas the remaining two were better represented by a double Sérsic model. Additionally, they find that a model consisting of two Sérsic profiles for the inner regions and an exponential profile for the outer region generally provides a good fit for all galaxies, which is their primary model for analysis. Following theoretical predictions, they conclude that the inner Sérsic profile represents a component formed by in-situ processes. The middle Sérsic profile represents the dominant component composed of a phase-mixed accreted component where the material has dynamically relaxed. The outer exponential profile was then associated with the unrelaxed accreted component, which often shows low surface brightness signatures of interactions, like shells or streams. From the 1D surface brightness profiles, they locate inflections that are indicative of physical transitions between the different components. A brief comparison with our 1D surface brightness profiles revealed that the inflection occurs at a galactocentric distance of kpc for the three ellipticals. The sizes of the inner “in-situ” component from our models lie within this apparent transition radius with kpc. However, given the different model combinations used in both studies and the fact that our 1D surface brightness profiles are constructed solely from the band image slice of the datacube, we cannot conclusively determine if the inner component corresponds to the transition radius observed in the surface brightness profiles, as this radius can shift depending on the wavelength employed.
Despite the distinct methodologies, both the photometric and spectro-photometric models support the idea that the inner component is primarily shaped by in-situ star formation, while the outer component is dominated by ex-situ accretion, consistent with theoretical predictions and simulations.
@@ 5.2 Stellar population properties: inside-out formation and the two-phase scenario
Several studies in literature have investigated the mode of mass assembly in elliptical galaxies using IFS. However, most of these studies address this question by measuring the radial gradients of stellar age and metallicity. Our methodology instead measures the mean stellar population properties in each of the two “clean” components extracted with the spectro-photometric code BUDDI, where the mixing of light is mitigated between the components. Therefore, the comparisons that we make to literature in this Section are less specific to the measurements and techniques, and rather focus on the general results and implications to elliptical galaxy formation.
@@
With the diverse range of methods and observations used to tackle this, there exists a broad variation in the stellar population results reported by different IFS-based studies. For instance, González Delgado et al. (2015) report elliptical galaxies in the Calar Alto Legacy Integral Field Area Survey (CALIFA) survey showing slight negative gradients in light-weighted age, with Gyr in the centre and dropping to Gyr at 1. A notable outcome by Zibetti et al. (2020) in their light-weighted stellar age analysis for 48 elliptical galaxies in CALIFA was the U-shaped profile for the ages, with the transition occurring at . Otherwise, in contrast to the other studies, they find a slightly positive gradient up to , beyond which the slope flattened out. The authors attribute this inversion to the interplay between competing quenching mechanisms: the outside-in quenching due to feedback-driven winds is favoured at large radial distances, while inside-out quenching triggered by AGN-driven feedback occurs starting from the galactic centre moving outwards. Moreover, they suggest that these results lend support to the two-phase formation scenario, wherein the quenching mechanisms likely compete during the initial phase. The flattening of the slope at the outermost radial regions could then be attributed to the accretion of ex-situ stars from satellite galaxies, where their effect is the most prominent. With the Mapping Nearby Galaxies at APO (MaNGA) survey, Lacerna et al. (2020) studied the stellar population properties of classical ellipticals, recently quenched ellipticals, and blue star-forming ellipticals independently. Our sample of ellipticals fall under the classical category, and we limit our comparisons to this sub-sample. They found moderately negative gradients in the light-weighted ages, and flat gradients in the mass-weighted ages. Similar flat age gradients were reported in Parikh et al. (2019), while weak negative age gradients were found in Domínguez Sánchez et al. (2019) for elliptical galaxies in MaNGA, and similarly in Riffel et al. (2023) for ETGs with MaNGA. A different study with MaNGA by Goddard et al. (2017) reported contrasting slightly positive gradients for ETGs, which include both ellipticals and S0s.
@@
In principle, from these studies, a positive gradient in stellar age implies the galaxy has assembled its mass from the outside in. A negative gradient, on the other hand, suggests an inside-out mass growth, starting from the innermost regions of the present-day observed elliptical galaxy. Along those lines, the flat gradients observed in some works would indicate a more uniform mass growth throughout the galaxy. However, the stellar populations of galaxies, like most of their properties, are driven by a multitude of processes that are more likely to be interwoven rather than relying on a single mechanism. Our results from the stellar population analysis fall within the wide diversity of results that have been observed in the studies mentioned above. Both the mean mass-weighted and light-weighted stellar age differences between the inner and outer components were low to moderate ( Gyr). For the three elliptical galaxies in this sample, this observation insinuates a mild inside-out formation scenario for J020536 and J225546, and a more pronounced inside-out formation for J205050. This is qualitatively and globally similar to the moderately negative or flat gradients reported in a few of the aforementioned IFS-based studies. In the context of a two-component galaxy, the inner component would have formed stars earlier, while the star formation continued for a slightly longer period, building up the stellar mass on the outskirts. Such old stellar populations are typical for galaxies that have an early intense star formation in the inner component occurring in an extremely short timescale, which quench soon after a few Gyr. The relatively younger ages in the outer component could then be associated to accreted stars from satellite galaxies that are themselves quenched and are nearly as old as the main elliptical galaxy. At this point, we note again the high levels of degeneracy associated with stellar populations older than 8 Gyr, and it is important to be cautious when interpreting such old ages.
@@
While the mean stellar ages offer some perspective into the growth mode of these galaxies, they are solely not enough to confirm this theory. The list of literature results described earlier in this Section additionally report results on the stellar metallicity analyses for their samples of ellipticals and ETGs. Most of the work done along these lines in the past with IFS have consistently reported negative metallicity gradients, with only the slope signifying their steepness varying between them. The exception is González Delgado et al. (2015), where they find nearly flat gradients in mass-weighted and light-weighted stellar metallicities for an elliptical galaxy sample in CALIFA. The reason for this discrepancy is not fully understood, and could be linked to the different methodologies employed in different studies. For an extensively shared sample, Zibetti et al. (2020) on the other hand, notice steeper gradients up to , after which the profiles transition to a flatter slope. Consistent with this, for their sample of classical ellipticals, Lacerna et al. (2020) find mildly steep gradients for the mass-weighted and light-weighted metallicities up to , beyond which the profiles flatten. Negative metallicity gradients are in accordance with the results presented Parikh et al. (2019), Domínguez Sánchez et al. (2019), and Goddard et al. (2017).
@@
Based on these studies on stellar metallicity gradients, a negative slope points to one possible scenario of early dissipative collapse from a gas-rich cloud already enhanced with metals. The formation and metallicity profiles of ellipticals were investigated via simulations in Kobayashi (2004), where the relevant physical processes including radiative cooling, stellar feedback through SN and winds, and chemical enrichment models were incorporated. However, the AGN feedback often observed in massive galaxies was not considered. Under this simulated regime, they find that the radial metallicity gradients trace SFHs, and require a combination of both rapid dissipative collapse and major mergers to adequately explain their variation. Moreover, they expect that subsequent major mergers tend to taper out the radial metallicity profiles, diminishing signatures of physical processes prior to the merger. If the progenitor mass ratios are high enough, the metallicity gradients can be completely destroyed and would appear flat in present-day observations. Additionally, the variation in metallicity gradients can be triggered by late gas accretion in the outer regions, where a slower phase of star formation is then induced. With the inclusion of AGN feedback in Taylor & Kobayashi (2017), they reinforce the flattening of metallicity gradients by major mergers in massive galaxies, but the quenching induced by AGN makes it challenging to regenerate the gradients through later in-situ central star formation.
@@
Our results show a substantial decrease in the metal enhancement for the outer component, compared to the metal-rich inner component, observed in both the mass-weighted and light-weighted properties. A clear negative radial gradient of metallicity is seen qualitatively in the 2D maps as well, with a decline from the central region of the galaxies. This points to an intense in-situ star formation phase in the inner component, that might have been influenced by major mergers. This is difficult to discern from our analysis since our estimates depict the mean stellar metallicity for each component, and therefore any information on gradients that might or might not exist within the inner component is lost. The decline in metallicity for the outer component, in tandem with the old stellar ages presented by these stellar populations, can be explained on the basis of accretion of ex-situ stars from old satellite galaxies that have formed their stars from metal-poor gas. In a dry merger, the ex-situ stars are deposited on the outskirts of the galaxy due to its lower binding energy compared to the in-situ formed stars (Spavone et al., 2017), causing accreted stars to dominate in the outer component. The combined stellar population properties presented in this work are therefore compatible with the two-phase scenario described in Sect. 1, with a preferred inside-out mass growth for the intermediate-mass ellipticals comprising our sample.

### v2

@@ 4.4.2 Star formation histories from mass fractions
The 2D SFHs built from the mass fractions in general illustrate a picture where the inner component of the elliptical galaxies are extremely old and have higher metallicities. These weights are concentrated in the upper right corner of the grid, with a minute spread in age ( Gyr), while there is a larger spread in metallicity for galaxies J020536 and J225546, between -0.5 and 0.3 dex. The outer component also shows a dominant old stellar population, with a wider spread in age compared to the inner component, and an even greater spread in metallicity, with stellar populations spanning ages Gyr and the entire metallicity range of the grid. These are indicative of minor episodes of continuous and slower stellar mass assembly that have occurred after the primary peak of star formation, wherein different stellar populations have been accreted.
@@
The 1D mass-weighted SFHs trace the cumulative mass assembly of each component through cosmic time. They complement the results from the 2D SFHs with a clear depiction of the inner component having assembled its stellar mass initially and swiftly within a very short timescale, forming the inner core of the galaxy. The peak mass assembly of the outer component exhibits a delay of several gigayears compared to the inner component, and has assembled its mass at a slower rate, over the course of a few gigayears, most likely growing its mass by continuous accretion during this time until attaining the present-day mass. The SFHs for the inner and outer components exhibit a flat slope from Gyr and Gyr ago, respectively, up to , marking their quiescence during this period. This clearly points to an inside-out mass growth scenario for local elliptical galaxies.
@@
Furthermore, the timescales and were measured from the cumulative mass fractions (Table 4). measures the look-back time when half of the present-day stellar mass had been assembled, while measures the same for 90% of the present-day mass, which is related to the onset of quenching in the galaxy (Ferré-Mateu et al., 2018; Romero-Gómez et al., 2024), especially for retired quiescent galaxies. For all three ellipticals, the inner component had already formed half of their mass by Gyr ago, while the outer component had taken Gyr longer to assemble the same (with timescales of 2.09 Gyr, 4.74 Gyr, and 670 Myr for J020536, J205050, and J225546, respectively). Similarly, the inner component appears to have quenched its star formation activity as early as Gyr ago. The outer component, on the other hand, stopped forming stars only Gyr ago, roughly Gyr after half the stellar mass had assembled. The relative contribution of the inner component to the total stellar mass was found to be 50% for all three galaxies, indicating that half of the stellar mass originates from the outer component. These collective findings demonstrate an early and rapid assembly of the majority of the stellar content within the inner component of the elliptical galaxies, which consequently appears to be a ‘co-dominant’ component in terms of its influence on the present day total stellar mass. The longer assembly timescales attributed to the outer component, along with it hosting half of the galaxy stellar mass, imply a slower addition of stars well after the inner component has quenched its star formation.
@@ 4.4.4 Star formation histories from light fractions
The SFHs drawn from the light fractions are shown in the lower panels of Figs. 7, 8, and 9, respectively. The 1D cumulative light-weighted age distributions (right-most panels of the above-mentioned figures) serve as a qualitative proxy for the SFH, reflecting the relative contributions of stars formed at different cosmic epochs, to the current galaxy luminosity. Since the light-weighting is performed with the band luminosity, the results also trace when star formation episodes had occurred during more recent times. and derived from the cumulative light fractions (Table 4) indicate the look-back time when the stars contributed to 50% and 90%, respectively, of the luminosity of the galaxy or its components. In this case, does not necessarily highlight the onset of quiescence, but rather highlights the epoch when the stars contributed to 90% of the present-day luminosity. We find that the light-weighted 2D SFHs follow the trends observed in the mass-weighted SFHs, revealing a similar pattern in the history of star formation. In J205050, a significant portion of the light originates from stellar populations older than 5 Gyr, spanning a broad range of metallicities from super-solar to very sub-solar, reflecting a diverse range of stellar populations, similar to the mass-weighted 2D SFH. In contrast, the outer component of J225546 is dominated primarily by stars older than 10 Gyr, but with a wide spread in metallicities.
@@
The light fractions of J205050 similarly show a predominantly old and metal-enhanced light-weighted stellar population in the inner component, which, similar to J020536, has undergone a rapid event of bulk star formation at early epochs. The look-back time, indicating the major star formation in this component, seems to have occurred and ceased roughly 13.5 Gyr ago, followed by a flat unperturbed quiescent slope leading to the present-day. The outer component, however, shows a delayed contribution to 90% of the luminosity, with most of these stars having formed around 7 Gyr ago. It has remained quiescent since then, with no significant star formation up to the present day.
@@
Following this trend, the 2D light-weighted SFH of J225546 shows an inner component again dominated by the light of the old metal-rich stars at solar and super-solar metal abundances. The outer component consists of two distinct light-weighted stellar populations, one with old and metal-enhanced stars, and the other with similarly old but quite metal-poor stars reaching [M/H] dex. The cumulative light fractions complement this with a clear indication that the peak of star formation in both the inner and outer components happened at earlier look-back times. The outer component followed after the principal star formation occurred, but the formation timescale between both components for this galaxy is not as substantial as for the other two galaxies, which is also evident from their , indicating a delay of roughly 3 Gyr.
@@
The stellar population analyses and the reconstructed SFHs support the two-phase scenario introduced in Sect. 1. The inner regions of the ellipticals ( kpc) could, in principle, be associated with the in situ star formation within the galaxy as a result of dissipative collapse or major mergers very early on in their lifetimes. In contrast, the outer component ( kpc in this sample), is likely a combination of star formation formed in situ from gas, with the majority of the stellar envelope having been accreted through dry mergers over longer timescales compared to the inner component. Regarding the former, Choi et al. (2024) find that in cosmological simulations, mergers can result in the accretion of gas from external galaxies. They also find signatures of smoothly accreted gas, particularly through the cooling of the galaxy halo, aligning with the findings of Lagos et al. (2015) in ETGs. In Choi et al. (2024), another origin is attributed to gas from the primary galaxy halo that has been recycled over generations of stellar evolution. While these processes were examined in the context of gas accretion in the inner regions of galaxies fuelling a supermassive black hole, a smaller fraction of the gas can be distributed over different regions, including the outskirts. On the other hand, the accretion of direct stellar material from dry mergers with external satellites of different metallicities occurs around the old, metal-rich core, which undergoes minimal evolution since its star formation was quenched early on.
@@ 5.1 Multiple structural components in ellipticals: Implications from surface brightness profiles
The notion that elliptical galaxies may consist of multiple components has been examined in several studies through 2D photometric decomposition, which are outlined in this section. This work builds upon existing methods by incorporating the imaging and spectroscopic capabilities of IFS. The spectro-photometric decomposition technique with BUDDI not only models the surface brightness profiles of various components at each wavelength of an IFU datacube, but also enables a more accurate derivation of stellar populations from the resulting spectra compared to photometric colours, which often suffer from the age-metallicity degeneracy (Worthey, 1999).
@@
Huang et al. (2013a) investigated the possibility of multi-component ellipticals in a sample of 94 local galaxies in the CGS survey. The high spatial resolution of deep optical images from this survey allowed for a reliable 2D decomposition of the ellipticals into three to four sub-components using GALFIT (Peng et al., 2002, 2011), so that they could study their structural properties in the band. They found that for the majority of their sample the surface brightness profiles were best modelled by three Sérsic components, representing a compact inner core, a middle component, and an outer extended envelope. The 2D residuals consistently improved with increasing complexity in the models. Interestingly, they reported only a small fraction of their sample to be modelled with two components. Furthermore, they also suggested that their high-luminosity sample of ellipticals showed photometric signatures of being ‘core’ ellipticals with a shallow slope within the innermost component. Lacerna et al. (2016) report similar findings, noting a good fit with three Sérsic models for a sample of 89 isolated local elliptical galaxies from SDSS images. Huang et al. (2013a) and Huang et al. (2013b) linked the inner and intermediate components to the phase of elliptical galaxy formation involved in situ dissipative processes, while associating the outer component with a later stage of ex situ star accretion through dry mergers. These findings and interpretations closely align with the ones presented in this work, despite differences in the number of components modelled.
@@
Along these lines, Spavone et al. (2017) performed a detailed analysis of the surface brightness profiles of six massive ETGs in A VST Early-type GAlaxy Survey (VEGAS) survey (Capaccioli et al., 2015). They adopted various profile combinations to identify the most accurate model. Four galaxies were effectively modelled using a Sérsic + exponential profile, whereas the remaining two were better represented by a double Sérsic model. Additionally, they found that a model consisting of two Sérsic profiles for the inner regions and an exponential profile for the outer region generally provides a good fit for all galaxies; this model is the primary one they used for the analysis. Following theoretical predictions, they conclude that the inner Sérsic profile represents a component formed by in situ processes. The middle Sérsic profile represents the dominant component composed of a phase-mixed accreted component where the material has dynamically relaxed. The outer exponential profile is then associated with the unrelaxed accreted component, which often shows low-surface-brightness signatures of interactions, like shells or streams. From the 1D surface brightness profiles, they located inflections that are indicative of physical transitions between the different components. A brief comparison with our 1D surface brightness profiles reveals that the inflection occurs at a galactocentric distance of kpc for the three ellipticals. The sizes of the inner in situ component from our models lie within this apparent transition radius with kpc. However, given the different model combinations used in both studies and the fact that our 1D surface brightness profiles are constructed solely from the band image slice of the datacube, we cannot conclusively determine if the inner component corresponds to the transition radius observed in the surface brightness profiles, as this radius can shift depending on the wavelength employed. Despite the distinct methodologies, both the photometric and spectro-photometric models support the idea that the inner component is primarily shaped by in situ star formation, while the outer component is dominated by ex situ accretion, which is consistent with theoretical predictions and simulations.
@@ 5.2 Stellar population properties: Inside-out formation and the two-phase scenario
Several studies in the literature have investigated the mode of mass assembly in elliptical galaxies using IFS. However, most of these studies address this question by measuring the radial gradients of stellar age and metallicity. Our methodology instead measures the mean stellar population properties in each of the two ‘clean’ components extracted with the spectro-photometric code BUDDI, where the mixing of light is mitigated between the components. Therefore, the comparisons that we make to literature in this section are less specific to the measurements and techniques, and rather focus on the general results and implications to elliptical galaxy formation.
@@
With the diverse range of methods and observations used to tackle this, there exists a broad variation in the stellar population results reported by different IFS-based studies. For instance, González Delgado et al. (2015) report elliptical galaxies in the Calar Alto Legacy Integral Field Area Survey (CALIFA) survey showing slight negative gradients in light-weighted age, with Gyr in the centre and dropping to Gyr at 1. A notable outcome by Zibetti et al. (2020) in their light-weighted stellar age analysis for 48 elliptical galaxies in CALIFA was the U-shaped profile for the ages, with the transition occurring at . Otherwise, in contrast to the other studies, they find a slightly positive gradient up to , beyond which the slope flattened out. The authors attribute this inversion to the interplay between competing quenching mechanisms: the outside-in quenching due to feedback-driven winds is favoured at large radial distances, while inside-out quenching triggered by AGN-driven feedback occurs starting from the galactic centre moving outwards. Moreover, they suggest that these results lend support to the two-phase formation scenario, wherein the quenching mechanisms likely compete during the initial phase. The flattening of the slope at the outermost radial regions could then be attributed to the accretion of ex situ stars from satellite galaxies, where their effect is the most prominent. With the Mapping Nearby Galaxies at APO (MaNGA) survey, Lacerna et al. (2020) studied the stellar population properties of classical ellipticals, recently quenched ellipticals, and blue star-forming ellipticals independently. Our sample of ellipticals fall under the classical category, and we limit our comparisons to this sub-sample. They found moderately negative gradients in the light-weighted ages, and flat gradients in the mass-weighted ages. Similar flat age gradients were reported in Parikh et al. (2019), while weak negative age gradients were found in Domínguez Sánchez et al. (2019) for elliptical galaxies in MaNGA, and similarly in Riffel et al. (2023) for ETGs with MaNGA. A different study with MaNGA by Goddard et al. (2017) reported contrasting slightly positive gradients for ETGs, which include both ellipticals and S0s.
@@
In principle, from these studies, a positive gradient in stellar age implies the galaxy has assembled its mass from the outside in. A negative gradient, on the other hand, suggests an inside-out mass growth, starting from the innermost regions of the present-day observed elliptical galaxy. Along those lines, the flat gradients observed in some works would indicate a more uniform mass growth throughout the galaxy. However, the stellar populations of galaxies, like most of their properties, are driven by a multitude of processes that are more likely to be interwoven rather than relying on a single mechanism. Our results from the stellar population analysis fall within the wide diversity of results that have been observed in the studies mentioned above. Both the mean mass-weighted and light-weighted stellar age differences between the inner and outer components were low to moderate ( Gyr). For the three elliptical galaxies in this sample, this observation insinuates a mild inside-out formation scenario for J020536 and J225546, and a more pronounced inside-out formation for J205050. This is qualitatively and globally similar to the moderately negative or flat gradients reported in a few of the aforementioned IFS-based studies. In the context of a two-component galaxy, the inner component would have formed stars earlier, while the star formation continued for a slightly longer period, building up the stellar mass on the outskirts. Such old stellar populations are typical for galaxies that have an early intense star formation in the inner component occurring in an extremely short timescale, which quench soon after a few gigayears. The younger ages in the outer component could then be associated with accreted stars from satellite galaxies that are themselves quenched and are nearly as old as the main elliptical galaxy. At this point, we note again the high levels of degeneracy associated with stellar populations older than 8 Gyr, and it is important to be cautious when interpreting such old ages.
@@
While the mean stellar ages offer some perspective into the growth mode of these galaxies, they are solely not enough to confirm this theory. The list of literature results described earlier in this section additionally report results on the stellar metallicity analyses for their samples of ellipticals and ETGs. Most of the work done along these lines in the past with IFS have consistently reported negative metallicity gradients, with only the slope signifying their steepness varying between them. The exception is González Delgado et al. (2015), where they find nearly flat gradients in mass-weighted and light-weighted stellar metallicities for an elliptical galaxy sample in CALIFA. The reason for this discrepancy is not fully understood, and could be linked to the different methodologies employed in different studies. For an extensively shared sample, Zibetti et al. (2020), on the other hand, notice steeper gradients up to , after which the profiles transition to a flatter slope. Consistent with this, for their sample of classical ellipticals, Lacerna et al. (2020) find mildly steep gradients for the mass-weighted and light-weighted metallicities up to , beyond which the profiles flatten. Negative metallicity gradients are in accordance with the results presented Parikh et al. (2019), Domínguez Sánchez et al. (2019), and Goddard et al. (2017).
@@
Based on these studies on stellar metallicity gradients, a negative slope points to one possible scenario of early dissipative collapse from a gas-rich cloud already enhanced with metals. The formation and metallicity profiles of ellipticals were investigated via simulations in Kobayashi (2004), where the relevant physical processes including radiative cooling, stellar feedback through SN and winds, and chemical enrichment models were incorporated. However, the AGN feedback often observed in massive galaxies was not considered. Under this simulated regime, they find that the radial metallicity gradients trace SFHs, and require a combination of both rapid dissipative collapse and major mergers to adequately explain their variation. Moreover, they expect that subsequent major mergers tend to taper out the radial metallicity profiles, diminishing signatures of physical processes prior to the merger. If the progenitor mass ratios are high enough, the metallicity gradients can be completely destroyed and would appear flat in present-day observations. Additionally, the variation in metallicity gradients can be triggered by late gas accretion in the outer regions, where a slower phase of star formation is then induced. With the inclusion of AGN feedback in Taylor & Kobayashi (2017), they reinforce the flattening of metallicity gradients by major mergers in massive galaxies, but the quenching induced by AGN makes it challenging to regenerate the gradients through later in situ central star formation.
@@
Our results show a substantial decrease in the metal enhancement for the outer component, compared to the metal-rich inner component, observed in both the mass-weighted and light-weighted properties. A clear negative radial gradient of metallicity is seen qualitatively in the 2D maps as well, with a decline from the central region of the galaxies. This points to an intense in situ star formation phase in the inner component, which might have been influenced by major mergers. This is difficult to discern from our analysis since our estimates depict the mean stellar metallicity for each component, and therefore any information on gradients that might or might not exist within the inner component is lost. The decline in metallicity for the outer component, in tandem with the old stellar ages presented by these stellar populations, can be explained on the basis of accretion of ex situ stars from old satellite galaxies that have formed their stars from metal-poor gas. In a dry merger, the ex situ stars are deposited on the outskirts of the galaxy due to its lower binding energy compared to the in situ formed stars (Spavone et al., 2017), causing accreted stars to dominate in the outer component. The combined stellar population properties presented in this work are therefore compatible with the two-phase scenario described in Sect. 1, with a preferred inside-out mass growth for the intermediate-mass ellipticals comprising our sample.